OpenStudy (anonymous):

Factor by grouping. 15u^3+10u^2-105u

OpenStudy (anonymous):

= 5u*3u^2 + 5u*2u - 5u*21 = 5u(3u^2 + 2u - 21) = 5u(3u^2 + 9u - 7u - 21) = 5u[3u(u + 3) -7(u + 3)] = 5u[(3u - 7)(u + 3)] = 5u(3u - 7)(u + 3)